y^3tdy+1/2y^4dt=0


1
Expert's answer
2021-06-10T02:32:53-0400
"y^3tdy+\dfrac{1}{2}y^4 dt=0"

"y=0\ or \ \dfrac{dy}{y}=-\dfrac{1}{2t}dt"

"\int \dfrac{dy}{y}=\int(-\dfrac{1}{2t})dt"

"\ln|y|=-\dfrac{1}{2}\ln|t|+\ln c"

"y=\dfrac{c\sqrt{x}}{x}\ or\ y=0"